# Translating Flow to Policy via Hindsight Online Imitation

### v1

@@ 1 Introduction
The success of vision and language models is deeply rooted in their ability to train on vast quantities of data (brown2020language; ravi2024sam). However, in robotics, this reliance on large datasets presents a key challenge for traditional end-to-end approaches like imitation learning (lin2024data). Consequently, hierarchical robotic systems have garnered considerable interest as a way to circumvent this fundamental data bottleneck. These systems structure robotic control problems into two hierarchical levels, where a high-level planner decomposes abstract task specifications into subgoals to guide the low-level controller. This decomposition separates the high-level planner from the robot’s complex physical dynamics, thereby enabling it to be trained on vast, action-free datasets, even including non-robot data.
Flow of points is a general representation of these high-level plans, describing the predicted state as future keypoint trajectories (wen2023ATManypointtrajectory; xu2025im2flow2act; haldar2025pointpolicy). Since point flow explicitly encodes physical motion dynamics and is robust to variations in visual appearance, it offers compact and effective guidance for the low-level controller. However, translating these flow plans into robust and scalable low-level policies remains a critical challenge. A direct approach is using analytical or optimization-based methods to compute actions, but they often struggle with real-world complexities, e.g., visual occlusions or non-rigid-body dynamics (bharadhwaj2024track2act; yuan2024general). Alternatively, learning a data-driven low-level policy is a more general method; however, it typically relies on collecting high-quality, in-domain robot data (wen2023ATManypointtrajectory; gao2024flip), which is an expensive and laborious process that undermines the goal of scalable skill acquisition.
To address these challenges, we propose Hindsight Flow-conditioned Online Imitation (HinFlow), a simple yet effective approach that enables the robot to ground the high-level flows into executable policies through self-practice. The robot interacts with the environment under the flow guidance to generate exploratory behaviors and continuously refine a flow-conditioned control policy. The core insight is that even if the collected experiences fail to reach the planner’s precise goal, they can be repurposed for self-imitation by framing the achieved flows as the intended goal (ghoshlearning). Generating such supervision directly from the robot’s own imperfect experiences allows it to learn and adapt without relying on a large volume of expert data, thus constructing a scalable bridge from flow plans to robust physical execution.
We demonstrate the effectiveness of HinFlow through experiments on benchmark tasks from LIBERO (liu2023libero) and ManiSkill (taomaniskill3). Our key contributions are summarized as follows:
@@ 2.1 Hierarchical Robotic Systems
Hierarchical robotic systems structure control problems into a high-level planner, which generates subgoal guidance, and a low-level controller tasked with achieving them. Previous works have explored various goal representations, such as future images (du2023learning; blackzero; yang2023learning), point flows (wen2023ATManypointtrajectory; bharadhwaj2024towards), motion fields (yin2025object), or learned neural representations (wang2023mimicplay; fang2023generalization). Despite this diversity in goal formats, methods for mapping them to low-level control mainly fall into a few categories. One approach uses analytic or optimization-based controllers that, for example, compute an optimal rigid-body transform to align the current state with the goal (kolearning; yuan2024general; yin2025object). While direct, such methods inherently rely on assumptions like object rigidity. Another thread is a data-driven approach, which learns policy models for low-level control. Although general, prior work typically trains a goal-conditioned policy or inverse dynamics model on high-quality, in-domain robot data (wang2023mimicplay; du2023learning; wen2023ATManypointtrajectory), a process that requires substantial human effort in data collection and limits scalability. Consistent with our approach, recent work also considers involving online interactions to improve policy. For example, escontrela2023video uses high-level video prediction to construct rewards and trains policies via reinforcement learning (RL). However, RL with visual rewards typically suffers from inefficient exploration and the complexities of optimizing long-horizon rewards (liu2023imitation). In contrast, we frame policy learning as supervised learning on hindsight-relabeled examples. This yields a simple, well-conditioned objective that is easy to optimize. Moreover, zhou2025autonomous also employs hindsight relabeling to improve goal-conditioned policies. Unlike their image-goal formulation, we leverage point flows as a high-level plan, which is a compact, low-dimensional encoding of task-relevant motion. This abstraction filters out non‑task‑relevant visual variations, leading to a more robust control policy.
@@ 2.2 Flow-based Manipulation
Point-motion (flow) extraction from visual inputs offers a general feature representation. Leveraging this representation, prior work has successfully guided manipulation tasks using point flows derived from diverse sources. Some approaches train flow prediction models on action-free or in-the-wild video datasets (wen2023ATManypointtrajectory; yuan2024general; bharadhwaj2024track2act). An alternative stream facilitates learning from human demonstrations by adopting object-centric flows (xu2025im2flow2act; wang2025skil) or unifying keypoint representations between human hands and robot grippers (haldar2025pointpolicy; ren2025motion). gao2024flip treats flow as an action representation within a world model, performing model-based planning to derive the optimal flow. gu2023rt further explores using manually drawn or LLM-generated trajectory sketches to provide effective guidance.
Once the flows are obtained, the subsequent flow-to-policy translation mirrors the paradigms outlined in Section 2.1. Our approach is closely related to research on flow-based reinforcement learning. guzey2025hudorbridging leverages object point flows in a human video, defining a reward that encourages the agent to replicate these motion trajectories. Similarly, yu2025genflowrl uses a flow derived from a generative model to shape an object-centric dense reward function. However, unlike prior methods that require predicting the full, long-horizon flow (point motions across an entire rollout), we utilize short-horizon flows as the high-level goal. While this difference might appear small, it has two major implications: (1) It is a significant challenge for the high-level model to predict the long-horizon flow. (2) Furthermore, using short-horizon flows allows us to reframe the problem as self-imitation. In our framework, the policy learns to replicate the achieved flow sequences from collected experiences, offering an intuitive and effective policy improvement algorithm.
@@ 3 Preliminary
Policy Learning from Videos and Demonstrations. The action-free video dataset is denoted by $\mathcal{D}_{h}=\{\tau^{i}_{o}\}_{i=1}^{N_{h}}$, where $N_{h}$ denotes the number of videos. Following existing research (wen2023ATManypointtrajectory), $\mathcal{D}_{h}$ is leveraged to train a high-level planner that produces subgoals $\mathcal{G}_{t}$ at each timestep $t$. In addition, we assume access to an action-labeled dataset $\mathcal{D}_{a}=\{\tau^{i}_{a}\}_{i=1}^{N_{a}}$, where each trajectory $\tau^{i}_{a}=\{(o_{t}^{i},a_{t}^{i})\}_{t=1}^{T}$ consists of observation–action pairs of length $T$. These demonstrations are typically employed to train a low-level imitation policy $\pi$ to execute the task under the guidance of the high-level planner. Formally, the policy parameters $\theta$ are optimized as follows:
@@
Point Flow. We adopt point flow as our high-level plan representation. Point flow captures motion patterns from videos while being invariant to task-irrelevant factors like appearance and lighting. To extract point flow from videos, we first select a set of initial points $\mathbf{p}_{0}=\{p_{0,k}\}_{k=1}^{K}$ in the first camera frame, where $p_{0,k}=(x_{0,k},y_{0,k})$ denotes pixel coordinates. We then apply an off-the-shelf video tracker (karaev2024cotracker) to obtain their 2D trajectories $\mathbf{p}_{t}$ for $t=1,\dots,T$. These trajectories constitute the ground-truth labels for training the high-level planner. The trained planner predicts the subgoal at time $t$ in the format of $\mathcal{G}_{t}=\{\hat{\mathbf{p}}_{i}\}_{i=t}^{t+H}$, where $H$ denotes the planning horizon.
@@ 4.1 High-level Planner with Flow Modeling from Videos
To realize the flow-based planner, we first extract flow information from the video dataset $\mathbf{D}_{h}$ for the high-level training process. As detailed in Section 3, we use an off-the-shelf video tracker $\Phi$ (karaev2024cotracker3) to generate flow labels for the video dataset. Specifically, for each observation $o_{t}$ in $\mathcal{D}_{h}$, we apply our task-centric point sampler to generate a set of query points $\mathbf{p}_{t}$ and then track their future trajectory $\mathbf{p}^{t+H}_{t+1}$. Such data tuple $(o_{t},\mathbf{p}^{t+H}_{t})$ are stored into the annotated video dataset $\mathcal{\bar{D}}_{h}$. Further data augmentation details are provided in Appendix A.1.
After obtaining the flow annotations, we proceed to train the flow prediction model. Following the Track Transformer model in ATM (wen2023ATManypointtrajectory), we tokenize the images and query points, and feed them into a multi-modal transformer model. The model is trained using a flow prediction loss, which can be formally expressed as:
@@ 4.2 Low-level Policy with Hindsight Online Imitation
To collect data for policy training, the robot interacts with the environment using its hierarchical policy. To facilitate exploration and enable the policy to encounter novel states, we introduce a modest amount of exploration noise into the action, as detailed in Appendix A.1. The environment responds to the output action by providing a new observation and state. This process is iterated until the episode terminates. After recording one full episode $\{(o_{t},a_{t})\}_{t=1}^{T}$, the pipeline conducts hindsight relabeling (kaelbling1993learning; andrychowicz2017hindsight; ghoshlearning; yang2022rethinking) to produce valid training data. To annotate these data with flow labels, we utilize the video tracker $\Phi$ to process the collected data, generating the achieved flow. The point sampling strategy and data augmentation are consistent with those described in Section 4.1. The labeled data $(o_{t},a_{t},\{\mathbf{p}_{i}\}_{i=t}^{t+H})$ for $t=1,\dots,T$ are subsequently incorporated into the replay buffer $\mathcal{D}_{r}$.
@@
The architecture of our flow-conditioned low-level policy adheres to the transformer-based design established in prior research (kim2021vilt; wen2023ATManypointtrajectory). Specifically, the inputs of the current timestep, comprising both visual and proprioceptive data, are initially encoded into a transformer. These input modalities are subsequently encoded into spatial tokens through a spatial transformer. We then combined flow tokens with the spatial tokens to effectively integrate the guiding information from the tracks. The combined tokens are then passed through a series of MLPs to generate actions. Furthermore, we incorporate action chunking into the policy learning process, as elaborated in Appendix A.1.
@@ 5.1 Experiment Setup
To rigorously assess HinFlow’s capabilities, we conduct experiments on seven tasks from two benchmarks. The suite includes four tasks from LIBERO (liu2023libero): Place Butter, Place Book, Hide Chocolate, Close Microwave, and three from ManiSkill3 (taomaniskill3): Place Sphere, Pull Cube Tool, and Poke Cube. This collection presents a diverse set of challenges that span long-horizon manipulation (e.g., Hide Chocolate) and complex object interaction (e.g., Pull Cube Tool). Figure 3 illustrates the initial and target state for each task.
@@
  • ATM (wen2023ATManypointtrajectory) first pretrains a track transformer on large-scale video dataset to predict the future flow of the query points. The official implementation queries a set of points on a fixed grid, which we denote as ATM (grid). We further implement a variant named ATM (seg), which is equipped with our point sampling strategy that concentrates on the robot end-effector and task-relevant objects.
  • Online VPT adapts the original Video PreTraining (VPT) (baker2022video) to our setting. VPT first trains an inverse dynamics model (IDM) on action-labeled demonstrations. The trained IDM is then used to provide action labels for a large-scale video dataset, from which a BC policy is learned. Our primary modification is an iterative online refinement loop: it uses online rollouts from the policy to fine-tune the IDM, then uses the improved IDM to relabel the dataset and retrain the policy.
@@ 5.3 Real World Experiment
As shown in Figure 5, the task requires a Franka Emika Panda arm with a Robotiq gripper to grasp and place a mouse onto a pad. The robot operates at a control frequency of 10 Hz, with an action space consisting of the end-effector translations and the gripper command. We employ a wrist-mounted ZED2 camera and a fixed third-person ZED2 camera to capture the images. The RGB images are then cropped and resized to $\times128$ resolution as observations. We leverage Grounded SAM 2 (ravi2024sam; ren2024grounded) and CoTracker3 (karaev2024cotracker3) to extract point flows. For each episode, the mouse is manually reset to a randomized position within a 15cm $\times$ 15cm area (red box in Figure 5).
Data collection is performed via VR teleoperation using a Meta Quest controller, following the hardware and system setup established in DROID (khazatsky2024droid). A dataset of 100 action-free videos is used for training the high-level planner, and only 2 action-labeled expert demonstrations are used to pretrain the low-level policy.
@@ 6 Conclusion, Limitations and Future Work
Despite its effectiveness, our work has several limitations. First, to ensure collecting meaningful rollouts in the early exploration phase, we initialize our low-level policy with a small number of in-domain, action-labeled demonstrations. Recent research on vision-language-action models with open-world capabilities (zitkovich2023rt; kim2024openvla; black2024pi_0) could potentially provide these initial exploratory trajectories, thereby eliminating the need for any action-labeled data. Second, our work is built upon 2D point flow, which is inherently ambiguous for complex 3D motions. Therefore, extending our method to more advanced goal representations, such as 3D motion fields (yin2025object), presents a valuable direction for future investigation.

### v2

@@ 1 Introduction
The success of vision and language models is deeply rooted in their ability to train on vast quantities of data (Brown et al., 2020; Ravi et al., 2024). However, in robotics, this reliance on large datasets presents a key challenge for traditional end-to-end approaches like imitation learning (Lin et al., 2024). Consequently, hierarchical robotic systems have garnered considerable interest as a way to circumvent this fundamental data bottleneck. These systems structure robotic control problems into two hierarchical levels, where a high-level planner decomposes abstract task specifications into subgoals to guide the low-level controller. This decomposition separates the high-level planner from the robot’s complex physical dynamics, thereby enabling it to be trained on vast, action-free datasets, even including non-robot data.
Flow of points is a general representation of these high-level plans, describing the predicted state as future keypoint trajectories (Wen et al., 2023; Xu et al., 2025; Haldar and Pinto, 2025). Since point flow explicitly encodes physical motion dynamics and is robust to variations in visual appearance, it offers compact and effective guidance for the low-level controller. However, translating these flow plans into robust and scalable low-level policies remains a critical challenge. A direct approach is using analytical or optimization-based methods to compute actions, but they often struggle with real-world complexities, e.g., visual occlusions or non-rigid-body dynamics (Bharadhwaj et al., 2024b; Yuan et al., 2024). Alternatively, learning a data-driven low-level policy is a more general method; however, it typically relies on collecting high-quality, in-domain robot data (Wen et al., 2023; Gao et al., 2024), which is an expensive and laborious process that undermines the goal of scalable skill acquisition.
To address these challenges, we propose Hindsight Flow-conditioned Online Imitation (HinFlow), a simple yet effective approach that enables the robot to ground the high-level flows into executable policies through self-practice. The robot interacts with the environment under the flow guidance to generate exploratory behaviors and continuously refine a flow-conditioned control policy. The core insight is that even if the collected experiences fail to reach the planner’s precise goal, they can be repurposed for self-imitation by framing the achieved flows as the intended goal (Ghosh et al., 2019). Generating such supervision directly from the robot’s own imperfect experiences allows it to learn and adapt without relying on a large volume of expert data, thus constructing a scalable bridge from flow plans to robust physical execution.
We demonstrate the effectiveness of HinFlow through experiments on benchmark tasks from LIBERO (Liu et al., 2023) and ManiSkill (Tao et al., 2025). Our key contributions are summarized as follows:
@@ 2.1 Hierarchical Robotic Systems
Hierarchical robotic systems structure control problems into a high-level planner, which generates subgoal guidance, and a low-level controller tasked with achieving them. Previous works have explored various goal representations, such as future images (Du et al., 2023; Black et al., 2023; Yang et al., 2023), point flows (Wen et al., 2023; Bharadhwaj et al., 2024a), motion fields (Yin et al., 2025), or learned neural representations (Wang et al., 2023; Fang et al., 2023). Despite this diversity in goal formats, methods for mapping them to low-level control mainly fall into a few categories. One approach uses analytic or optimization-based controllers that, for example, compute an optimal rigid-body transform to align the current state with the goal (Ko et al., 2023; Yuan et al., 2024; Yin et al., 2025). While direct, such methods inherently rely on assumptions like object rigidity. Another thread is a data-driven approach, which learns policy models for low-level control. Although general, prior work typically trains a goal-conditioned policy or inverse dynamics model on high-quality, in-domain robot data (Wang et al., 2023; Du et al., 2023; Wen et al., 2023), a process that requires substantial human effort in data collection and limits scalability. Consistent with our approach, recent work also considers involving online interactions to improve policy. For example, Escontrela et al. (2023) uses high-level video prediction to construct rewards and trains policies via reinforcement learning (RL). However, RL with visual rewards typically suffers from inefficient exploration and the complexities of optimizing long-horizon rewards (Liu et al., 2024). In contrast, we frame policy learning as supervised learning on hindsight-relabeled examples. This yields a simple, well-conditioned objective that is easy to optimize. Moreover, Luo and Du and Zhou et al. (2025) both utilize hindsight relabeling to enhance goal-conditioned policies: Luo and Du employs video diffusion models to guide online exploration, while Zhou et al. (2025) leverages vision-language models to generate task goal proposals. Unlike their image-goal formulation, we leverage point flows as a high-level plan, which is a compact, low-dimensional encoding of task-relevant motion. This abstraction filters out non‑task‑relevant visual variations, leading to a more robust control policy.
@@ 2.2 Flow-based Manipulation
Point-motion (flow) extraction from visual inputs offers a general feature representation. Leveraging this representation, prior work has successfully guided manipulation tasks using point flows derived from diverse sources. Some approaches train flow prediction models on action-free or in-the-wild video datasets (Wen et al., 2023; Yuan et al., 2024; Bharadhwaj et al., 2024b). An alternative stream facilitates learning from human demonstrations by adopting object-centric flows (Xu et al., 2025; Wang et al., 2025) or unifying keypoint representations between human hands and robot grippers (Haldar and Pinto, 2025; Ren et al., 2025). Gao et al. (2024) treats flow as an action representation within a world model, performing model-based planning to derive the optimal flow. Gu et al. (2023) further explores using manually drawn or LLM-generated trajectory sketches to provide effective guidance.
Once the flows are obtained, the subsequent flow-to-policy translation mirrors the paradigms outlined in Section 2.1. Our approach is closely related to research on flow-based reinforcement learning. Guzey et al. (2025) leverages object point flows in a human video, defining a reward that encourages the agent to replicate these motion trajectories. Similarly, Yu et al. (2025) uses a flow derived from a generative model to shape an object-centric dense reward function. However, unlike prior methods that require predicting the full, long-horizon flow (point motions across an entire rollout), we utilize short-horizon flows as the high-level goal. While this difference might appear small, it has two major implications: (1) It is a significant challenge for the high-level model to predict the long-horizon flow. (2) Furthermore, using short-horizon flows allows us to reframe the problem as self-imitation. In our framework, the policy learns to replicate the achieved flow sequences from collected experiences, offering an intuitive and effective policy improvement algorithm.
@@ 3 Preliminary
Policy Learning from Videos and Demonstrations. The action-free video dataset is denoted by $\mathcal{D}_{h}=\{\tau^{i}_{o}\}_{i=1}^{N_{h}}$, where $N_{h}$ denotes the number of videos. Following existing research (Wen et al., 2023), $\mathcal{D}_{h}$ is leveraged to train a high-level planner that produces subgoals $\mathcal{G}_{t}$ at each timestep $t$. In addition, we assume access to an action-labeled dataset $\mathcal{D}_{a}=\{\tau^{i}_{a}\}_{i=1}^{N_{a}}$, where each trajectory $\tau^{i}_{a}=\{(o_{t}^{i},a_{t}^{i})\}_{t=1}^{T}$ consists of observation–action pairs of length $T$. These demonstrations are typically employed to train a low-level imitation policy $\pi$ to execute the task under the guidance of the high-level planner. Formally, the policy parameters $\theta$ are optimized as follows:
@@
Point Flow. We adopt point flow as our high-level plan representation. Point flow captures motion patterns from videos while being invariant to task-irrelevant factors like appearance and lighting. To extract point flow from videos, we first select a set of initial points $\mathbf{p}_{0}=\{p_{0,k}\}_{k=1}^{K}$ in the first camera frame, where $p_{0,k}=(x_{0,k},y_{0,k})$ denotes pixel coordinates. We then apply an off-the-shelf video tracker (Karaev et al., 2024b) to obtain their 2D trajectories $\mathbf{p}_{t}$ for $t=1,\dots,T$. These trajectories constitute the ground-truth labels for training the high-level planner. The trained planner predicts the subgoal at time $t$ in the format of $\mathcal{G}_{t}=\{\hat{\mathbf{p}}_{i}\}_{i=t}^{t+H}$, where $H$ denotes the planning horizon.
@@ 4.1 High-level Planner with Flow Modeling from Videos
To realize the flow-based planner, we first extract flow information from the video dataset $\mathbf{D}_{h}$ for the high-level training process. As detailed in Section 3, we use an off-the-shelf video tracker $\Phi$ (Karaev et al., 2024a) to generate flow labels for the video dataset. Specifically, for each observation $o_{t}$ in $\mathcal{D}_{h}$, we apply our task-centric point sampler to generate a set of query points $\mathbf{p}_{t}$ and then track their future trajectory $\mathbf{p}^{t+H}_{t+1}$. Such data tuple $(o_{t},\mathbf{p}^{t+H}_{t})$ are stored into the annotated video dataset $\mathcal{\bar{D}}_{h}$. Further data augmentation details are provided in Appendix A.1.
After obtaining the flow annotations, we proceed to train the flow prediction model. Following the Track Transformer model in ATM (Wen et al., 2023), we tokenize the images and query points, and feed them into a multi-modal transformer model. The model is trained using a flow prediction loss, which can be formally expressed as:
@@ 4.2 Low-level Policy with Hindsight Online Imitation
To collect data for policy training, the robot interacts with the environment using its hierarchical policy. To facilitate exploration and enable the policy to encounter novel states, we introduce a modest amount of exploration noise into the action, as detailed in Appendix A.1. The environment responds to the output action by providing a new observation and state. This process is iterated until the episode terminates. After recording one full episode $\{(o_{t},a_{t})\}_{t=1}^{T}$, the pipeline conducts hindsight relabeling (Kaelbling, 1993; Andrychowicz et al., 2017; Ghosh et al., 2019; Yang et al., 2022) to produce valid training data. To annotate these data with flow labels, we utilize the video tracker $\Phi$ to process the collected data, generating the achieved flow. The point sampling strategy and data augmentation are consistent with those described in Section 4.1. The labeled data $(o_{t},a_{t},\{\mathbf{p}_{i}\}_{i=t}^{t+H})$ for $t=1,\dots,T$ are subsequently incorporated into the replay buffer $\mathcal{D}_{r}$.
@@
The architecture of our flow-conditioned low-level policy adheres to the transformer-based design established in prior research (Kim et al., 2021; Wen et al., 2023). Specifically, the inputs of the current timestep, comprising both visual and proprioceptive data, are initially encoded into a transformer. These input modalities are subsequently encoded into spatial tokens through a spatial transformer. We then combined flow tokens with the spatial tokens to effectively integrate the guiding information from the tracks. The combined tokens are then passed through a series of MLPs to generate actions. Furthermore, we incorporate action chunking into the policy learning process, as elaborated in Appendix A.1.
@@ 5.1 Experiment Setup
To rigorously assess HinFlow’s capabilities, we conduct experiments on seven tasks from two benchmarks. The suite includes four tasks from LIBERO (Liu et al., 2023): Place Butter, Place Book, Hide Chocolate, Close Microwave, and three from ManiSkill3 (Tao et al., 2025): Place Sphere, Pull Cube Tool, and Poke Cube. This collection presents a diverse set of challenges that span long-horizon manipulation (e.g., Hide Chocolate) and complex object interaction (e.g., Pull Cube Tool). Figure 3 illustrates the initial and target state for each task.
@@
  • ATM (Wen et al., 2023) first pretrains a track transformer on large-scale video dataset to predict the future flow of the query points. The official implementation queries a set of points on a fixed grid, which we denote as ATM (grid). We further implement a variant named ATM (seg), which is equipped with our point sampling strategy that concentrates on the robot end-effector and task-relevant objects.
  • Online VPT adapts the original Video PreTraining (VPT) (Baker et al., 2022) to our setting. VPT first trains an inverse dynamics model (IDM) on action-labeled demonstrations. The trained IDM is then used to provide action labels for a large-scale video dataset, from which a BC policy is learned. Our primary modification is an iterative online refinement loop: it uses online rollouts from the policy to fine-tune the IDM, then uses the improved IDM to relabel the dataset and retrain the policy.
@@ 5.3 Real World Experiment
As shown in Figure 5, the task requires a Franka Emika Panda arm with a Robotiq gripper to grasp and place a mouse onto a pad. The robot operates at a control frequency of 10 Hz, with an action space consisting of the end-effector translations and the gripper command. We employ a wrist-mounted ZED2 camera and a fixed third-person ZED2 camera to capture the images. The RGB images are then cropped and resized to $\times128$ resolution as observations. We leverage Grounded SAM 2 (Ravi et al., 2024; Ren et al., 2024) and CoTracker3 (Karaev et al., 2024a) to extract point flows. For each episode, the mouse is manually reset to a randomized position within a 15cm $\times$ 15cm area (red box in Figure 5).
Data collection is performed via VR teleoperation using a Meta Quest controller, following the hardware and system setup established in DROID (Khazatsky et al., 2024). A dataset of 100 action-free videos is used for training the high-level planner, and only 2 action-labeled expert demonstrations are used to pretrain the low-level policy.
@@ 6 Conclusion, Limitations and Future Work
Despite its effectiveness, our work has several limitations. First, to ensure collecting meaningful rollouts in the early exploration phase, we initialize our low-level policy with a small number of in-domain, action-labeled demonstrations. Recent research on vision-language-action models with open-world capabilities (Zitkovich et al., 2023; Kim et al., 2024; Black et al., 2024) could potentially provide these initial exploratory trajectories, thereby eliminating the need for any action-labeled data. Second, our work is built upon 2D point flow, which is inherently ambiguous for complex 3D motions. Therefore, extending our method to more advanced goal representations, such as 3D motion fields (Yin et al., 2025), presents a valuable direction for future investigation.